# Taxonomy of reduction matrices for Graph Coarsening

### v1

@@ 1 Introduction
We start by background material on graph coarsening in Sec. 2, highlighting the roles of the lifting and reduction matrices. We emphasize the asymmetricity of their roles, along with the strong constraints put on the lifting matrix alone. Then in Sec. 3, we study sets of admissible reduction matrices, focusing on various notions of “generalized inverses”. Some will be very generic, while others will admit parametrizations that are convenient for optimization. In Sec. 4, we then study several motivated examples of reduction matrices, classical or entirely novel, with an analytical closed-form expression or based on optimization procedures. We relate them to the properties defined in the section before. Finally, we compare their performance in terms of RSA or GNN performance in Sec. 5. The proofs are deferred to App. B.
@@ 2 Characterizing graph coarsening with the lifting matrix
A graph $G$ with $N$ nodes is described by its weighted adjacency matrix $A\in\mathbb{R}^{N\times N}$. The combinatorial Laplacian is defined as $\mathcal{L}=\mathcal{L}(A):=D-A$, where $D=D(A):=\text{diag}(A1_{N})$ is the diagonal matrix of the degrees. A matrix is said to be binary if all its coefficients are either 00 or $1$. For a symmetric positive semi-definite (p.s.d) matrix $M$, we denote $\|x\|_{M}=\sqrt{x^{\top}Mx}$ the Mahalanobis semi-norm associated with $M$.
@@
In addition to be well-partitioned, it is also natural to impose that ${\mathcal{Q}}$ is binary, as shown by [26, Prop.7] in the following Lemma.
@@
Classically, the preserved subspace $\mathcal{R}$ is spanned by the eigenvectors of the first eigenvalues of $\mathcal{L}$. In this case, the RSA constant can be used to bound the deviation between the spectrums of $\mathcal{L}$ and $\mathcal{L}_{c}$. When $\mathcal{R}$ is a eigen-subspace of $\mathcal{L}$, the RSA has an explicit expression [26]: $\epsilon_{\mathcal{L},{\mathcal{Q}},P,\mathcal{R}}=\lVert\mathcal{L}^{1/2}(I_{N}-P{\mathcal{Q}})VV^{T}\mathcal{L}^{+1/2}\rVert_{2}$ where $V$ is an orthogonal basis of $\mathcal{R}$ and $\lVert\cdot\rVert_{2}$ is the spectral norm. Note that this expression is convex in $P$, which will be useful for optimization.
@@
To extend the definition of the RSA constant (3) to the case of the generalized ${\Delta}$-Laplacian, it is first necessary to ensure that the norms used in the original and coarsened graphs, $G$ and $G_{c}$, are comparable [26, Corollary 12]. This requires establishing the consistency of the ${\Delta}$-Laplacian matrices of $G$ and $G_{c}$. This is shown with the following lemma, provided that, starting from any binary well-partitioned lifting matrix ${\mathcal{Q}}\in\mathbb{R}^{N\times n}$, a generalized lifting matrix is constructed as
@@
where ${\Delta}_{c}={\Delta}(A_{c})$ with $A_{c}$ defined in (1). Note that ${Q}$ is also well-partitioned when ${\mathcal{Q}}$ is well-partitioned, however it is generally not binary. Instead, the constraint still lies on ${\mathcal{Q}}$, as shown below.
@@
    For all adjacency matrices $A$, we have $L(A_{c})={Q}(A,{\mathcal{Q}})^{\top}L(A){Q}(A,{\mathcal{Q}})$, where we recall that $L$ is defined in (4) and ${Q}$ in (5).
@@
where, again, we emphasize that $L$ is defined in (4) and ${Q}$ in (5). The main question that we will examine in the rest of the paper remains: for a fixed ${\mathcal{Q}}$ that is well-partitioned and binary, what are the “valid” reduction matrices $P$, and is there a more “optimal” choice to minimize $\epsilon_{L,Q,P,\mathcal{R}}$?
@@ 3 Expanding the space of reduction matrices
In the beginning of this section, $Q$ will indicate any well-partitioned matrix. We note that, unlike all the matrices considered in the literature, some matrices $P$ with a support different from $Q^{\top}$ will be acceptable. Intuitively, ${Q}$ defines the “true” mapping between the nodes and the supernodes by enforcing the well-partitioned aspect, while $P$ will be allowed to relax this constraint. We begin by examining the largest such ensemble, denoted $E_{1}$.
@@
    $AB$ is Hermitian: $\left(AB\right)^{*}=AB$.
@@
    $BA$ is Hermitian: $\left(BA\right)^{*}=BA$.
@@
    generalized inverse of $A$ when it satisfies  (i);
@@
    generalized reflexive inverse of $A$ if it satisfies simultaneously (i) and (ii);
@@
    Moore Penrose inverse of $A$ if it satisfies  (i), (ii), (iii) and (iv).
@@
Let $Q={Q}(A,{\mathcal{Q}})\in\mathbb{R}^{N\times n}$ be generalized lifting matrix with ${\mathcal{Q}}$ well-partitioned and binary. The set of reflexive generalized inverse of ${Q}$ with the same support as ${Q}^{\top}$ is defined as :
@@
Note that, while Lem. 3, 4 and 5 were valid for any well-partitioned matrix ${Q}$, here we specifically examine $Q={Q}(A,{\mathcal{Q}})$ with ${\mathcal{Q}}$ well-partitioned and binary. Moreover, all matrices with non-zero coefficients on the support of ${Q}^{\top}$ that are also in $E_{1}$ are in $E_{2}$ and $E_{3}$, as their rank is equal to $n$.
@@ 4 From classical to novel reduction matrices: a comparative study
Now that we have proposed a taxonomy of the valid reduction matrices $P$, we investigate the second question raised in Sec. 2: what are good examples of reduction matrices? In all this section, we consider $Q={Q}(A,{\mathcal{Q}})$ with ${\mathcal{Q}}$ well-partitioned and binary. As outlined in Sec. 2, $L_{c}=Q^{\top}LQ$ is then the $\Delta$-Laplacian of the coarsened graph. We will start with three examples with closed-form analytic expression, then outline a possible optimization framework, emphasizing that it is only one choice among many possible.
@@
As mentioned before, when $\mathcal{R}$ is a subspace of ${L}$, the RSA has a close form expression that is convex in $P$. This results in optimization problem that are convex in $M$ in these cases. In our experiments, we treat the $l_{1}$ penalty simply with gradient descent combined with a final thresholding operation (for simplicity we leave aside more complex optimization procedures with e.g. proximal operators).
@@ 5 Experiments
(ii) Laplacian and preserved space. We choose two different Laplacians: the combinatorial Laplacian $\mathcal{L}$ and the self-loop normalized Laplacian ${L}={\Delta}\mathcal{L}{\Delta}$ with ${\Delta}=(\operatorname{diag}(A1_{N})+1)^{-1/2}$. The motivation for this Laplacian comes from the fact that it is related to the propagation matrix $S=\hat{D}^{-1/2}(A+I_{N})\hat{D}^{-1/2}$ (via ${L}=I_{N}-S$) commonly used in GNNs [23, 22]. For the RSA, the preserved space $\mathcal{R}$ is chosen as the $K=100$ first eigenvectors of $\mathcal{L}$ and ${L}$.
(iii) Lifting matrix ${Q}$. The method that has introduced the notion of RSA [26] is particularly relevant for computing ${Q}$, as we share the same objective of reducing the RSA constant. However, it requires adaptation because this method minimizes the RSA constant for the combinatorial Laplacian, and yields a binary well-partitioned matrix lifting matrix ${\mathcal{Q}}$. We generalize this method to minimize the RSA constant for a $\Delta$-Laplacian, resulting in a lifting matrix $Q$. It is worth noting that some of the assumptions made in [26] no longer hold in the $\Delta$-Laplacian setting. Nevertheless, we observe that this generalized construction still achieves good RSA constants. The details of the generalized algorithm are provided in App. F.
@@
The RSA constants achieved by the reduction matrices introduced in Sec.  4 are shown in Fig. 2(a) for the combinatorial Laplacian $\mathcal{L}$ and in Fig. 2(b) for the self-loop normalized Laplacian ${L}$. The coarsening ratios range from 0.05 to 0.85. As expected, the two proposed methods ${P_{opt}}$ and $P^{*}_{g}$ achieve better RSA constants than the usual $P_{MP}$ and ${P_{Loukas}}$. Moreover, the performance gap increases with higher reduction ratios. Another interesting observation is that the optimization methods $P^{*}_{g,l_{1}}$ and $P^{*}_{Q^{\top}}$ that incorporate sparsity constraints also perform very well. At low reduction ratios, their performance is nearly indistinguishable from their unconstrained counterparts $P^{*}_{g}$. This is particularly surprising given that the reduction in the number of non-zero coefficients is around $99.8\%$ , regardless of the coarsening ratios (see App. G.2 for a detailed analysis on the sparsity of the $P$ matrices and App. E for their computational hyperparameters). We also observe that the two approaches $P^{*}_{g,l_{1}}$ and $P^{*}_{Q^{\top}}$ yield similar performance. This may be due to the strong regularization parameter used in the sparsity penalty in $P^{*}_{g,l_{1}}$; a different setting could yield a more balanced trade-off for $P^{*}_{g,l_{1}}$ between the performance of $P^{*}_{g}$ and $P^{*}_{Q^{\top}}$.
@@
(b) Cora graph, self-loop normalized Laplacian ${L}$
@@
Inspired by the paper [22] which link the training of Graph Neural Networks (GNN) and the RSA : we have trained for three different coarsening ratio ( $r=\{0.3,0.5,0.7\}$) a convolutional GNN [23] and a Simplified convolution network (SGC [37]) on Cora [29] and Citeseer [14]. Each training is averaged on $10$ random split, using the same experimental setting as in [22], and the hyperparameters are provided in App. G.1. For the training on coarsened graphs, we used the propagation matrix $S_{c}^{MP}=PSQ$ from the paper [22], where $S-L$ and $L$ is the self-loop normalized Laplacian. This matrix depends on both the reduction matrix $P$ and the lifting matrix $Q$.
@@
[…]
@@ 6 Conclusion
In this paper, we highlighted the crucial role of the lifting matrix ${Q}$ in graph coarsening. Surprisingly, we found that we can take advantage of the degree of freedom over the reduction matrix to obtain better spectral guarantees, without changing the coarsening itself or the lifting matrix. We have defined various sets of “admissible” reduction matrices with different properties, from the very generic property of simply obtaining a projection $\Pi$, to convenient parametrization with reflexive generalized inverses, and support constraints. We then showed that the classical choices of reduction matrices can be outperformed, both by well-motivated novel examples with analytic expressions, or by matrices resulting from various optimization processes with sparsity or support constraints. Even if previous works linked the performances of GNNs trained on coarsened graph with the RSA, we empirically showed that the benefits of improving the RSA were somewhat marginal, although visible for high coarsening ratio and homophilous graphs. This suggests that other factors, such the sparsity of the reduction matrix, could also play in GNNs training.
In this work, we have selected the RSA as a general-purpose score to optimize, however our theoretical characterization of the admissible sets of reduction matrices does not particularly rely on it. We thus believe that considering more complex scoring function to take into account graphs heterophily or node features while being scalable to larger graphs is a major path for future works.
Ackowledgments
The authors acknowledge the fundings of France 2030, PEPR IA, ANR-23-PEIA-0008 and the European Union ERC-2024-STG-101163069 MALAGA.
@@ Appendix A Useful property and definitions
Indeed if one term is different from 00, it means that two nodes $i$ and $j$ of $G_{c}$ have been expanded to a same node $k$ of $G$ which contradicts the well-partitioned definition.
@@ B.3 Proof ${P_{Loukas}}$ generalized reflexive inverse
We apply this characterization to the well-partitioned lifting matrix ${Q}\in\mathbb{R}^{N\times n}$ using a well known generalized inverse of ${Q}$, namely the "Moore-Penrose pseudo inverse" $Q^{+}$ that has been characterized in App. B.2.
@@
=$=\begin{cases}0\text{ when }i\neq j\\ 1\text{ when}i=j\end{cases}$

### v2

@@ 1 Introduction
We start by background material on graph coarsening in Sec. 2, highlighting the roles of the lifting and reduction matrices. We emphasize the asymmetricity of their roles, along with the strong constraints put on the lifting matrix alone. Then in Sec. 3, we study sets of admissible reduction matrices, focusing on various notions of “generalized inverses”. Some will be very generic, while others will admit parametrizations that are convenient for optimization. In Sec. 4, we then study several motivated examples of reduction matrices, classical or entirely novel, with an analytical closed-form expression or based on optimization procedures. We relate them to the properties defined in the section before. Finally, we compare their performance in terms of RSA or GNN performance in Sec. 5. The code is available at https://gitlab.inria.fr/anjoly/taxonomy-coarsening-matrices, and proofs are deferred to App. B.
@@ 2 Characterizing graph coarsening with the lifting matrix
A graph $G$ with $N$ nodes is described by its weighted adjacency matrix $A\in\mathbb{R}^{N\times N}$. The combinatorial Laplacian is defined as $\mathcal{L}=\mathcal{L}(A):=D-A$, where $D=D(A):=\text{diag}(A1_{N})$ is the diagonal matrix of the degrees. A matrix is said to be binary if all its coefficients are either 0 or $1$. For a symmetric positive semi-definite (p.s.d) matrix $M$, we denote $\|x\|_{M}=\sqrt{x^{\top}Mx}$ the Mahalanobis semi-norm associated with $M$.
@@
In addition to be well-partitioned, it is also natural to impose that $\mathcal{Q}$} is binary, as shown by [28, Prop.7] in the following Lemma.
@@
Classically, the preserved subspace $\mathcal{R}$ is spanned by the eigenvectors of the first eigenvalues of $\mathcal{L}$. In this case, the RSA constant can be used to bound the deviation between the spectrums of $\mathcal{L}$ and $\mathcal{L}_{c}$. When $\mathcal{R}$ is an eigen-subspace of $\mathcal{L}$, the RSA has an explicit expression [28]: $\epsilon_{\mathcal{L},{\mathcal{Q}},P,\mathcal{R}}=\lVert\mathcal{L}^{1/2}(I_{N}-P{\mathcal{Q}})VV^{T}\mathcal{L}^{+1/2}\rVert_{2}$ where $V$ is an orthogonal basis of $\mathcal{R}$ and $\lVert\cdot\rVert_{2}$ is the spectral norm. Note that this expression is convex in $P$, which will be useful for optimization. This definition slightly differs from [28], as it disentangles the roles of the lifting matrix $\mathcal{Q}$} and the reduction matrix $P$, while in [28], $\mathcal{Q}$} was fixed as the Moore Penrose inverse of $P$.
@@
To extend the definition of the RSA constant (3) to the case of the generalized $\Delta$}-Laplacian, it is first necessary to ensure that the norms used in the original and coarsened graphs, $G$ and $G_{c}$, are comparable [28, Corollary 12]. This requires establishing the consistency of the $\Delta$}-Laplacian matrices of $G$ and $G_{c}$. This is shown with the following lemma, provided that, starting from any binary well-partitioned lifting matrix ${\mathcal{Q}}\in\mathbb{R}^{N\times n}$, a generalized lifting matrix is constructed as
@@
where ${\Delta}_{c}={\Delta}(A_{c})$ with $A_{c}$ defined in (1). Note that Q{Q} is also well-partitioned when $\mathcal{Q}$} is well-partitioned, however it is generally not binary. Instead, the constraint still lies on $\mathcal{Q}$}, as shown below.
@@
    For all adjacency matrices $A$, we have $L(A_{c})={Q}(A,{\mathcal{Q}})^{\top}L(A){Q}(A,{\mathcal{Q}})$, where we recall that $L$ is defined in (4) and Q{Q} in (5).
@@
where, again, we emphasize that $L$ is defined in (4) and Q{Q} in (5). The main question that we will examine in the rest of the paper remains: for a fixed $\mathcal{Q}$} that is well-partitioned and binary, what are the “valid” reduction matrices $P$, and is there a more “optimal” choice to minimize $\epsilon_{L,Q,P,\mathcal{R}}$?
@@ 3 Expanding the space of reduction matrices
In the beginning of this section, $Q$ will indicate any well-partitioned matrix. We note that, unlike all the matrices considered in the literature, some matrices $P$ with a support different from $Q^{\top}$ will be acceptable. Intuitively, Q{Q} defines the “true” mapping between the nodes and the supernodes by enforcing the well-partitioned aspect, while $P$ will be allowed to relax this constraint. We begin by examining the largest such ensemble, denoted $E_{1}$.
@@
    ABAB is Hermitian: $\left(AB\right)^{*}=AB$.
@@
    BABA is Hermitian: $\left(BA\right)^{*}=BA$.
@@
    generalized inverse of AA when it satisfies  (i);
@@
    generalized reflexive inverse of AA if it satisfies simultaneously (i) and (ii);
@@
    Moore Penrose inverse of AA if it satisfies  (i), (ii), (iii) and (iv).
@@
Let $Q={Q}(A,{\mathcal{Q}})\in\mathbb{R}^{N\times n}$ be generalized lifting matrix with $\mathcal{Q}$} well-partitioned and binary. The set of reflexive generalized inverse of Q{Q} with the same support as ${Q}^{\top}$ is defined as :
@@
Note that, while Lem. 3, 4 and 5 were valid for any well-partitioned matrix Q{Q}, here we specifically examine $Q={Q}(A,{\mathcal{Q}})$ with $\mathcal{Q}$} well-partitioned and binary. Moreover, all matrices with non-zero coefficients on the support of ${Q}^{\top}$ that are also in $E_{1}$ are in $E_{2}$ and $E_{3}$, as their rank is equal to $n$. To provide a better intuition about Lem. 6, consider the case of the combinatorial Laplacian (${\Delta}=I_{N}$) where $Q={\mathcal{Q}}$; this lemma reduces to $\sum_{k=1}^{N}P_{ik}=1$, which appears to be a natural condition already used in [3].
@@ 4 From classical to novel reduction matrices: a comparative study
Now that we have proposed a taxonomy of the valid reduction matrices $P$, we investigate the second question raised in Sec. 2: what are good examples of reduction matrices? In all this section, we consider $Q={Q}(A,{\mathcal{Q}})$ with $\mathcal{Q}$} well-partitioned and binary. As outlined in Sec. 2, $L_{c}=Q^{\top}LQ$ is then the $\Delta$-Laplacian of the coarsened graph. We will start with three examples with closed-form analytic expression, then outline a possible optimization framework, emphasizing that it is only one choice among many possible. It is worth noting that the transposed matrix $P=Q^{\top}$, commonly used in graph pooling, does not belong to $E_{1}$, and is therefore not considered in our analysis.
@@
As mentioned before, when $\mathcal{R}$ is a subspace of L{L}, the RSA has a close form expression that is convex in $P$. This results in optimization problem that are convex in $M$ in these cases. In our experiments, we treat the $l_{1}$ penalty simply with gradient descent combined with a final thresholding operation (for simplicity we leave aside more complex optimization procedures with e.g. proximal operators).
@@ 5 Experiments
(ii) Laplacian and preserved space. We choose two different Laplacians: the combinatorial Laplacian $\mathcal{L}$ and the self-loop normalized Laplacian ${L}={\Delta}\mathcal{L}{\Delta}$ with ${\Delta}=(\operatorname{diag}(A1_{N})+1)^{-1/2}$. The motivation for this Laplacian comes from the fact that it is related to the propagation matrix $S=\hat{D}^{-1/2}(A+I_{N})\hat{D}^{-1/2}$ (via ${L}=I_{N}-S$) commonly used in GNNs [24, 23]. For the RSA, the preserved space $\mathcal{R}$ is chosen as the $K=100$ first eigenvectors of $\mathcal{L}$ and L{L}.
(iii) Lifting matrix Q{Q}. The method that has introduced the notion of RSA [28] is particularly relevant for computing Q{Q}, as we share the same objective of reducing the RSA constant. However, it requires adaptation because this method minimizes the RSA constant for the combinatorial Laplacian, and yields a binary well-partitioned matrix lifting matrix $\mathcal{Q}$}. We generalize this method to minimize the RSA constant for a $\Delta$-Laplacian, resulting in a lifting matrix $Q$. It is worth noting that some of the assumptions made in [28] no longer hold in the $\Delta$-Laplacian setting. Nevertheless, we observe that this generalized construction still achieves good RSA constants. The details of the generalized algorithm are provided in App. F.
@@
The RSA constants achieved by the reduction matrices introduced in Sec.  4 are shown in Fig. 2(a) for the combinatorial Laplacian $\mathcal{L}$ and in Fig. 2(b) for the self-loop normalized Laplacian L{L}. The coarsening ratios range from 0.05 to 0.85. As expected, the two proposed methods $P_{opt}$} and $P^{*}_{g}$ achieve better RSA constants than the usual $P_{MP}$ and $P_{Loukas}$}. Moreover, the performance gap increases with higher reduction ratios. Another interesting observation is that the optimization methods $P^{*}_{g,l_{1}}$ and $P^{*}_{Q^{\top}}$ that incorporate sparsity constraints also perform very well. At low reduction ratios, their performance is nearly indistinguishable from their unconstrained counterparts $P^{*}_{g}$. This is particularly surprising given that the reduction in the number of non-zero coefficients is around $99.8\%$ , regardless of the coarsening ratios (see App. G.2 for a detailed analysis on the sparsity of the $P$ matrices and App. E for their computational hyperparameters). We also observe that the two approaches $P^{*}_{g,l_{1}}$ and $P^{*}_{Q^{\top}}$ yield similar performance. This may be due to the strong regularization parameter used in the sparsity penalty in $P^{*}_{g,l_{1}}$; a different setting could yield a more balanced trade-off for $P^{*}_{g,l_{1}}$ between the performance of $P^{*}_{g}$ and $P^{*}_{Q^{\top}}$.
@@
(b) Cora graph, self-loop normalized Laplacian L{L}
@@
Inspired by the paper [23] which link the training of Graph Neural Networks (GNN) and the RSA : we have trained for three different coarsening ratio ( $r=\{0.3,0.5,0.7\}$) a convolutional GNN [24] and a Simplified convolution network (SGC [38]) on Cora [30] and Citeseer [15]. Each training is averaged on $10$ random split, using the same experimental setting as in [23], and the hyperparameters are provided in App. G.1. For the training on coarsened graphs, we used the propagation matrix $S_{c}^{MP}=PSQ$ from the paper [23], where SLS-L and $L$ is the self-loop normalized Laplacian. This matrix depends on both the reduction matrix $P$ and the lifting matrix $Q$.
@@
Table 1: Accuracy in $\%$ for node classification with SGC and GCNconv on different coarsening ratio
@@ 6 Conclusion
In this paper, we highlighted the crucial role of the lifting matrix Q{Q} in graph coarsening. Surprisingly, we found that we can take advantage of the degree of freedom over the reduction matrix to obtain better spectral guarantees, without changing the coarsening itself or the lifting matrix. We have defined various sets of “admissible” reduction matrices with different properties, from the very generic property of simply obtaining a projection $\Pi$, to convenient parametrization with reflexive generalized inverses, and support constraints. We then showed that the classical choices of reduction matrices can be outperformed, both by well-motivated novel examples with analytic expressions, or by matrices resulting from various optimization processes with sparsity or support constraints. Even if previous works linked the performances of GNNs trained on coarsened graph with the RSA, we empirically showed that the benefits of improving the RSA were somewhat marginal, although visible for high coarsening ratio and homophilous graphs. This suggests that other factors, such the sparsity of the reduction matrix, could also play in GNNs training.
In this work, we have selected the RSA as a general-purpose score to optimize, however our theoretical characterization of the admissible sets of reduction matrices does not particularly rely on it. We thus believe that considering more complex scoring function to take into account graphs heterophily or node features while being scalable to larger graphs is a major path for future works. Notably, our optimization framework is agnostic to the specific choice of scoring function and coarsening algorithm and directly applies to these extensions.
Acknowledgments and Disclosure of Funding
The authors acknowledge the fundings of France 2030, PEPR IA, ANR-23-PEIA-0008 and European Union ERC-2024-STG-101163069 MALAGA.
@@ Appendix A Useful property and definitions
Indeed if one term is different from 0, it means that two nodes $i$ and $j$ of $G_{c}$ have been expanded to a same node $k$ of $G$ which contradicts the well-partitioned definition.
@@ B.3 Proof $P_{Loukas}$} generalized reflexive inverse
We apply this characterization to the well-partitioned lifting matrix ${Q}\in\mathbb{R}^{N\times n}$ using a well known generalized inverse of Q{Q}, namely the "Moore-Penrose pseudo inverse" $Q^{+}$ that has been characterized in App. B.2.
@@
$=\begin{cases}0\text{ when }i\neq j\\ 1\text{ when}i=j\end{cases}$